Find the slope if possible, of the line passing through each pair of points (2,-1) and (6,3)
Answer by checkley77(12844)

(2,-1) and (6,3)
slope=(y2-y1)/(x2-x1)
slope=(3+1)/(6-2)
slope=4/4=1 ans. for the slope.